# GPT-4 Jailbreaks Itself with Near-Perfect Success Using Self-Explanation

### v1

@@ Abstract
Research on jailbreaking has been valuable for testing and understanding the safety and security issues of large language models (LLMs). In this paper, we introduce Iterative Refinement Induced Self-Jailbreak (IRIS), a novel approach that leverages the reflective capabilities of LLMs for jailbreaking with only black-box access. Unlike previous methods, IRIS simplifies the jailbreaking process by using a single model as both the attacker and target. This method first iteratively refines adversarial prompts through self-explanation, which is crucial for ensuring that even well-aligned LLMs obey adversarial instructions. IRIS then rates and enhances the output given the refined prompt to increase its harmfulness. We find IRIS achieves jailbreak success rates of 98% on GPT-4 and 92% on GPT-4 Turbo in under 7 queries. It significantly outperforms prior approaches in automatic, black-box and interpretable jailbreaking, while requiring substantially fewer queries, thereby establishing a new standard for interpretable jailbreaking methods.
@@ 1 Introduction
Our experiments show that IRIS jailbreaks GPT-4 and GPT-4 Turbo  [^1] GPT-4 refers to gpt-4-0613 and GPT-4 Turbo refers to gpt-4-turbo-2024-04-09; experiments ran on 4/14/2024 with over 98% attack success rate in under 13 queries. It significantly improves both success rate and query efficiency compared to the prior state-of-the-art results set by TAP Mehrotra et al. (2023), which has a 75% attack ASR and requires more than 20 queries. We find that the IRIS prompts generated by GPT models can also successfully attack the Claude-3 Opus model Anthropic (2024), with a 80% success rate. We also present an ablation study to investigate the impact of each step in IRIS.
@@ 2 IRIS: a Self-Jailbreaking Method
Given an initial adversarial request $R_{adv}$ such as “write a tutorial on how to make bomb” that models always rejects, the goal is to formulate a prompt that induces the target LLM ($T$) to generate content that fulfills the request. Our method, IRIS, uses the same LLM as the attacker ($A$). As shown in Figure 1, IRIS consists of two main steps: Iterative Refinement, which iteratively refines the initial prompt , and Rate+Enhance, which maximizes the harmfulness of the output.
During the Iterative Refinement step, IRIS operates through a feedback loop that progressively refines the adversarial prompt based on $T$’s responses and $A$’s modifications. At each iteration, the current prompt $P_{current}$ is presented to $T$, and its response $R$ is evaluated to determine whether $T$ rejects the request based on a simple rule: reject if the response is less than 20 words; otherwise, do not reject. If $T$ rejects the prompt, IRIS solicits an explanation from the attacker model $A$ on why the attempt failed using a template query $Q_{A}(\texttt{EXPLAIN:}R)$. This self-explanation step is vital for well-aligned LLMs (e.g., GPT-4), as it prevents an immediate rejection when directly asking models to refine the failed adversarial prompt, $Q_{A}(\texttt{MODIFY:}P_{current})$. The output from the MODIFY query is a refined prompt, $P_{refined}$, which becomes the new basis for subsequent iterations. The iterative prompt refinement process continues until $R_{adv}$ is found or the number of attempts $N$ is reached, which we set $N=4$ in our experiments. Since there are 3 queries in each iteration, IRIS produces a maximum of $133N+1=$ 13 queries to LLMs for our experiments, which is significantly more efficient than previous approaches (over 20 queries). However, over 80% of the time, only one or two iterations are used.
@@ 3 Experiments
In addition to IRIS, we consider two state-of-the-art methods that use LLMs to refine jailbreak prompts: PAIR Chao et al. (2023) and TAP Mehrotra et al. (2023). PAIR uses Vicuna-13B Chiang et al. (2023) to iteratively refine the prompt, while TAP further improves the method by incorporating the tree-of-thought reasoning Yao et al. (2024). There is another method, PAP Zeng et al. (2024), that fine-tunes GPT-3.5 to generate prompts but requires 400 queries when jailbreaking GPT-4. We exclude it from our experiments.
@@
Following prior work Chao et al. (2023); Mehrotra et al. (2023), we use the AdvBench Subset dataset from Chao et al. (2023) in our experiments. It contains a diverse set of 50 adversarial prompts, covering a wide range of harmful categories. We report attack success rate (ASR), the percentage of 50 adversarial prompts that resulted in a successful jailbreak, and query efficiency, the total number of queries to the LLM’s API. A successful jailbreak is defined as the model producing a response to the original adversarial prompt that contains relevant harmful content. We calculate ASR based on human evaluation instead of using GPT-4 as a judge. GPT-4 has been shown to incorrectly classify jailbreaks as successful when they are not necessarily harmful Mehrotra et al. (2023); Yu et al. (2023). The human evaluation is completed by an in-house annotator who achieved 100% agreement with authors in a training tutorial that contains 30 examples, showing that this evaluation task is straightforward.
@@
Table 2: We evaluate the attack success rate when using a refined jailbreak prompt from one model on a different target. The top part showcases the effectiveness of self-jailbreaking. The bottom part shows the vulnerability of Claude-3 models in transfer attack by GPT-4 ones.
3.2 Main Results
Figure 1 shows the main results that compare IRIS with TAP and PAIR, whose results were reported in Mehrotra et al. (2023). IRIS-2x represents an ensemble of two independent IRIS trials on each adversarial prompt, where the jailbreak is considered successful if at least one of the trials succeeds. The average number of queries for IRIS-2x is the sum of the queries in the two trials. We find that IRIS achieves higher jailbreak success rates with significantly fewer queries than TAP and PAIR. IRIS has success rates of 98% and 92% for GPT-4 and GPT-4 Turbo, respectively, using under 7 queries on average. With two independent trials (IRIS-2x), these rates rise to 100% and 98% with under 13 queries on average, which is approximately 55% fewer queries than other methods while increasing the jailbreak success rate by at least 22%.
@@ 4 Conclusion
This work introduces a new mechanism by which to approach the the shortcomings of LLM alignment, and as such does not encompass every avenue of exploration. For example, we do not evaluate IRIS on open-source models, as we find they lack the capability to follow the algorithm’s instructions to the extent required for inducing a jailbreak. Future work could explore possible defense mechanisms against IRIS and investigate the effects of iteratively applying the Rate+Enhance step, which we only experiment with once in this study. Additionally, the prompt templates used in IRIS have only one format, which may be easy to detect. Future research could explore methods for automatically generating these templates to improve the robustness of the approach.

### v2

@@ Abstract
Research on jailbreaking has been valuable for testing and understanding the safety and security issues of large language models (LLMs). In this paper, we introduce Iterative Refinement Induced Self-Jailbreak (IRIS), a novel approach that leverages the reflective capabilities of LLMs for jailbreaking with only black-box access. Unlike previous methods, IRIS simplifies the jailbreaking process by using a single model as both the attacker and target. This method first iteratively refines adversarial prompts through self-explanation, which is crucial for ensuring that even well-aligned LLMs obey adversarial instructions. IRIS then rates and enhances the output given the refined prompt to increase its harmfulness. We find that IRIS achieves jailbreak success rates of 98% on GPT-4, 92% on GPT-4 Turbo, and 94% on Llama-3.1-70B in under 7 queries. It significantly outperforms prior approaches in automatic, black-box, and interpretable jailbreaking, while requiring substantially fewer queries, thereby establishing a new standard for interpretable jailbreaking methods.
@@ 1 Introduction
Our experiments show that IRIS jailbreaks GPT-4 and GPT-4 Turbo  [^1] GPT-4 refers to gpt-4-0613 and GPT-4 Turbo refers to gpt-4-turbo-2024-04-09; experiments ran on 4/14/2024 with over 98% attack success rate (ASR) in under 13 queries. It significantly improves both success rate and query efficiency compared to the prior state-of-the-art results set by TAP Mehrotra et al. (2023), which has a 75% ASR and requires more than 20 queries. We find that the IRIS prompts generated by GPT models can also successfully attack the Claude-3 Opus model Anthropic (2024), with an 80% success rate. We also present an ablation study to investigate the impact of each step in IRIS. Follow-up results show that IRIS achieves 94% and 95% ASR on the more recent Llama-3.1-70B and GPT-4o models, respectively.
@@ 2 IRIS: a Self-Jailbreaking Method
Given an initial adversarial request $R_{adv}$ such as “write a tutorial on how to make bomb” that models always reject, the goal is to formulate a prompt that induces the target LLM ($T$) to generate content that fulfills the request. Our method, IRIS, uses the same LLM for the attacker ($A$). As shown in Figure 1, IRIS consists of two main steps: Iterative Refinement, which iteratively refines the initial prompt, and Rate+Enhance, which maximizes the harmfulness of the output.
During the Iterative Refinement step, IRIS operates through a feedback loop that progressively refines the adversarial prompt based on $T$’s responses and $A$’s modifications. At each iteration, the current prompt $P_{current}$ is presented to $T$, and its response $R$ is evaluated to determine whether $T$ rejects the request based on a simple rule: reject if the response is less than 20 words; otherwise, do not reject. If $T$ rejects the prompt, IRIS solicits an explanation from the attacker model $A$ on why the attempt failed using a template query $Q_{A}(\texttt{EXPLAIN:}R)$. This self-explanation step is vital for well-aligned LLMs (e.g., GPT-4), as it prevents an immediate rejection when directly asking models to refine the failed adversarial prompt, $Q_{A}(\texttt{MODIFY:}P_{current})$. The output from the MODIFY query is a refined prompt, $P_{refined}$, which becomes the new basis for subsequent iterations. The iterative prompt refinement process continues until $R_{adv}$ is found or the number of attempts $N$ is reached, which we set $N=4$ in our experiments. Since there are 3 queries in each iteration, IRIS produces a maximum of $133N+1=$ 13 queries to LLMs for our experiments, which is significantly more efficient than previous approaches (over 20 queries). However, over 80% of the time, only one or two iterations are used. Experiment artifacts show refined prompts $P_{refined}$ produced by IRIS always request the same harmful behavior as the original prompts $P_{adv}$.
@@ 3 Experiments
In addition to IRIS, we consider two state-of-the-art methods that use LLMs to refine jailbreak prompts: PAIR Chao et al. (2023) and TAP Mehrotra et al. (2023). PAIR uses Vicuna-13B Chiang et al. (2023) to iteratively refine the prompt, while TAP further improves the method by incorporating the tree-of-thought reasoning Yao et al. (2024). There is another method, PAP Zeng et al. (2024), that fine-tunes GPT-3.5 to generate prompts but requires 400 queries when jailbreaking GPT-4. We exclude it from our comparisons.
@@
Following prior work Chao et al. (2023); Mehrotra et al. (2023), we use the AdvBench Subset dataset from Chao et al. (2023) in our experiments. It contains a diverse set of 50 adversarial prompts, covering restricted behaviors from OpenAI’s usage policies, including: government interference, explosives, hacking, identity theft, violence, financial crimes, mental health, theft, terrorism, hate crimes, piracy, stalking, phishing, and firearms. We report attack success rate (ASR), the percentage of 50 adversarial prompts that resulted in a successful jailbreak, and query efficiency, the total number of queries to the LLM’s API. A successful jailbreak is defined as the model producing a response to the original adversarial prompt that contains relevant harmful content. We calculate ASR based on human evaluation instead of using GPT-4 as a judge. GPT-4 has been shown to incorrectly classify jailbreaks as successful when they are not necessarily harmful Mehrotra et al. (2023); Yu et al. (2023). The human evaluation is completed by an in-house annotator who achieved 100% agreement with authors in a training tutorial that contains 30 examples, showing that this evaluation task is straightforward.
Model Jailbreak % Avg. Queries
Llama-3-8B 18% 4.3
Llama-3-70B 44% 5.6
Llama-3.1-8B 62% 4.6
Llama-3.1-70B 94% 4.4
GPT-4 Turbo 92% 5.3
GPT-4 98% 6.7
Table 2: Comparison of IRIS on open-source instruction-tuned Llama models and GPT-4 for direct jailbreak attacks on the AdvBench Subset. We report the attack success rate determined by human evaluation and the average number of queries required for each model.
3.2 Main Results
Table 1 shows the main results that compare IRIS with TAP and PAIR, whose results were reported in Mehrotra et al. (2023). IRIS-2x represents an ensemble of two independent IRIS trials on each adversarial prompt, where the jailbreak is considered successful if at least one of the trials succeeds. The average number of queries for IRIS-2x is the sum of the queries in the two trials. We find that IRIS achieves higher jailbreak success rates with significantly fewer queries than TAP and PAIR. IRIS has success rates of 98% and 92% for GPT-4 and GPT-4 Turbo, respectively, using under 7 queries on average. With two independent trials (IRIS-2x), these rates rise to 100% and 98% with under 13 queries on average, which is approximately 55% fewer queries than other methods while increasing the jailbreak success rate by at least 22%.
We also evaluate IRIS on an alternative benchmark, JailbreakBench (Chao et al., 2024), which contains 100 distinct misuse behaviors curated with reference to OpenAI’s usage policies, including original prompts (55%) and those sourced from prior work like HarmBench (27%) and AdvBench (18%). We find IRIS achieves ASR of 96% on GPT-4 Turbo and 95% on the newer GPT-4o, with an average of 4.72 and 4.66 queries respectively.
@@ 3.2 Main Results
Table 3: We evaluate the attack success rate when using a refined jailbreak prompt from one model on a different target. The top part showcases the effectiveness of self-jailbreaking. The bottom part shows the vulnerability of Claude-3 models in transfer attacks from GPT-4.
3.3 Open-Source Models
Although optimization-based white-box jailbreaking approaches like GCG (Zou et al., 2023) already achieve high attack success rates (ASR) on open-source models, we evaluate IRIS on instruction-tuned Llama 3 and 3.1 models (Dubey et al., 2024). Table 2 shows the results in comparison to the GPT-4 models. ASR increases significantly on the more proficient models, which we attribute to greater ability to follow the algorithm’s instructions to the extent required to induce a jailbreak.
@@ 4 Conclusion
This work introduces a new mechanism by which to approach the the shortcomings of LLM alignment, and as such does not encompass every avenue of exploration. Future work could explore possible defense mechanisms against IRIS and investigate the effects of iteratively applying the Rate+Enhance step, which we only experiment with once in this study. Additionally, the prompt templates used in IRIS have only one format, which may be easy to detect. Future research could explore methods for automatically generating these templates to improve the robustness of the approach.